OpenStudy (anonymous):

If a tree 13m long casts a 7m shadow how long is a shadow cast by a 2m person

OpenStudy (anonymous):

Tree Height / Man Height = Tree Shadow / Man Shadow 13 / 2 = 7 / Man Shadow Man Shadow = ?

OpenStudy (anonymous):

So do i cross multiply or what?

OpenStudy (anonymous):

\[\frac{13}{7}=\frac{2}{x}\]\[13x=14 \]\[x=\frac{14}{13} \]

OpenStudy (anonymous):

Tree Height / Man Height = Tree Shadow / Man Shadow 13 / 2 = 7 / Man Shadow Man Shadow = ? Man Shadow = 7 / 6.5 = ?

OpenStudy (anonymous):

@dauspex Yes. Method 1: Tree Height / Man Height = Tree Shadow / Man Shadow Method 2: Tree Height / Tree Shadow = Man Height / Man Shadow